OpenStudy (anonymous):

11. Write an equation in point-slope form for the line through the given point with the given slope (8, 3); m = 6 (1 point) A.y + 3 = 6(x – 8) B.y – 3 = 6(x – 8) C.y – 3 = 6(x + 8) D.y + 3 = 6x + 8

OpenStudy (anonymous):

y-y1=m(x-x1) b m=6